OpenStudy (anonymous):

Thirty is 12% of what number? A. 360 B. 2.5 C. 250 D. 3.6

OpenStudy (anonymous):

30=12/100*x, solve for x...

OpenStudy (anonymous):

The answer should be 250...

OpenStudy (texaschic101):

percent formula : is/of = %/100 is = 30 % = 12 of = x (sub in a variable to solve for because we are looking for this) now we sub 30/x = 12/100 cross multiply 12x = 3000 x = 3000/12 x = 250